# Anti-rotors and algebra invariants

### v36

@@ 1. Introduction
As the means of understanding algebra structure, identification of the invariants of an algebra is a task of fundamental importance as well as an open-ended endeavor given that the algebra classification problem is apparently intractable [1, 2]. While classical algebra isomorphism invariants include things like an algebra’s dimension, types of ideals, and invariants stemming from composition series (the Jordan-Hölder Theorem), more recently they have derived from module categories, e.g., involving homological formalisms such as found in [3, 4, 5]. The invariants we present here belong to a different class.
@@ 2. The Skewer and the anti-rotor
In (2.1), $\text{Sym}_{n}$ is the linear space of all possible metric tensors relevant to $\mathbb{R}^{n}$ (inclusive of metrics that are not necessarily positive-definite and not necessarily nondegenerate), and $\delta\text{Rot}_{n}$ is the linear space of all possible infinitesimal rotation tensors relevant to $\mathbb{R}^{n}$. Equation (2.2) is interpreted as follows. An algebra element is understood to be a column vector $s=(x_{1},\dots,x_{n})^{T}$, we define $\mathbf{d}s\equiv(dx_{1},\dots,dx_{n})^{T}$, and $\mu$ is a real symmetric $(n\times n)$-matrix. On the right-hand-side of (2.2), $d$ represents the exterior derivative operation, thus acting on the 1-form $(f(s))^{T}\mu\,\mathbf{d}s=\big(\mu f(s)\big)\cdot\mathbf{d}s$. Consequently, the right-hand-side of (2.2) yields the infinitesimal rotation tensor associated with the vector field $f(s)$ with respect to the metric $\mu$ - or, equivalently, the infinitesimal rotation tensor of vector field $\mu f(s)$ with respect to the Euclidean metric.
@@
$\not{\text{S}}_{f}$ is a linear transformation mapping symmetric tensors to skew-symmetric tensors, so we refer to $\not{\text{S}}_{f}$ as a “Skewer” (the slash through “symmetric” S represents a skewer, as with a shish-kabob). Since we assume the standard basis, $\text{Sym}_{n}$ can be represented by a parametrized symmetric $(n\times n)$-matrix with $\frac{n(n+1)}{2}$ independent parameters, i.e., it is given as a symmetric matrix with a different real parameter in each of its upper triangular entries. Every possible metric tensor is given by some realization of the set of parameters (a realization being the result of any particular choice of $\frac{n(n+1)}{2}$ real values for the members of the set of parameters). Consequently, this parametrized symmetric matrix is mapped by a Skewer to a parametrized skew-symmetric matrix representing the resulting infinitesimal rotation tensors. Each entry not on the main diagonal of the latter matrix is a rational function of the $n$ components of the elements of algebra $A$, and the above $\frac{n(n+1)}{2}$ parameters appear as coefficients of constituent rational function terms which sum to the entry.
While the image of a Skewer is thereby complicated, the kernel of a Skewer is much less so because the components of the elements of $A$ do not appear. Furthermore, as indicated in Section 8, there is suggestive computational evidence that the kernel of the Skewer is independent of the choice of $f$ over the class of functions that are dependent on the algebra’s product operation (note that for $f$ not dependent on the algebra product, the kernel of the Skewer is simply $\text{Sym}_{n}$ itself). Notwithstanding its pleasing simplicity, if the resulting independence conjecture is false it will only mean that the collection of invariants proceeding from this program is larger than we anticipate, thereby increasing its power. Either way, it will be seen that analysis of the mapping of algebras to Skewer kernels reveals previously hidden aspects of algebra structure.
@@
We will henceforth adopt “anti-rotor” as a more descriptive term for the kernel of the Skewer, because each of the latter’s members annihilates the curl of vector field $f$. For definiteness, in this paper we will work with $f(s)=s^{-1}$, since this choice is associated with a productive “dual” of the anti-rotor.
@@ 3. The anti-rotor and resulting invariants
where we have used Definition 3.1, the commutativity of a differential with a linear transformation, and $(Ks_{1})^{-1}=K(s_{1}^{-1})$ (since $K$ is an isomorphism). It follows from (3.3) that every uncurling metric of algebra $A_{2}$ is associated with an uncurling metric of $A_{1}$, i.e., from the last equality in (3.3) we have $K^{T}L_{2}K\in\mathfrak{u}_{A_{1}}$. Injectivity of the implied map from $\mathfrak{u}_{A_{2}}$ to $\mathfrak{u}_{A_{1}}$ is evident due to the invertibility of $K$. The isomorphism hypothesis allows us to apply an analogous argument employing $K^{-1}$, indicating that for $L_{1}\in\mathfrak{u}_{A_{1}}$ we have the associated $(K^{-1})^{T}L_{1}K^{-1}\in\mathfrak{u}_{A_{2}}$, implying an injective map from $\mathfrak{u}_{A_{1}}$ to $\mathfrak{u}_{A_{2}}$. The theorem then follows from the bijection implied by the Cantor-Schoeder-Bernstein Theorem. ∎
@@
For invertible elements $s=(s_{ij})$ in $M_{n}(\mathbb{R})$, with det(s)(s) being the determinant of $s$, we have the well-known,
@@
where Adj(s)(s) is the adjugate matrix, and the left-hand-side is considered to be a matrix according to $\nabla\text{det}(s)]\equiv\left(\frac{\partial\,\text{det}(s)}{\partial s_{ij}}\right)$. The first of these equalities is special case of Jacobi’s formula, and the second follows from the classical matrix inverse expression resulting from Laplace expansion. But equivalently, (3.4) can be understood as equating three vectors, where the matrix format of each bracketed expression is identified with the vector given by the $n^{2}$-tuple obtained by sequentially appending its $n$ columns, and the transpose operation is now understood as a particular involutive $(n^{2}\times n^{2})$-matrix, which we will denote as $\mathcal{T}$. In this vector format, after dividing both sides by $\text{det}(s)$, (3.4) becomes
@@
Now suppose the left regular representation of $\bar{A}$ is not self-adjoint as a subalgebra of $M_{n}(\mathbb{R})$. Then there exists an epimorphism $\Omega:\bar{A}\rightarrow\underline{A}$ such that $\underline{A}$ is a simple algebra, and therefore $\underline{A}$ has an uncurling metric $\underline{L}$. A simple modification of the first three sentences of the proof of Theorem 3.2 (where we exchange isomorphism $K$ for epimorphism $\Omega$), indicates that $\Omega^{T}\underline{L}\Omega$ satisfies (3.1) and is thus an uncurling metric for $\bar{A}$ - along with all its scalar multiples. That is, with respect to the elements $s_{2}\in\underline{A}$ and the elements $s_{1}\in\bar{A}$, we consider the linear function $s_{2}=\Omega s_{1}$. Then, $0=d\left((s_{2}^{-1})^{T}\underline{L}\mathbf{d}s_{2}\right)=d\left(((\Omega s_{1})^{-1})^{T}\underline{L}\mathbf{d}(\Omega s_{1})\right)=d\left((s_{1}^{-1})^{T}\big[\Omega^{T}\underline{L}\Omega\big]\mathbf{d}s_{1}\right)$, where we have used Definition 3.1, the commutativity of a differential with a linear transformation, and $(\Omega s_{1})^{-1}=\Omega(s_{1}^{-1})$ (since $\Omega$ is an epimorphism of rings). ∎
@@
The members of $\mathfrak{u}$ are the realizations of a symmetric $(n\times n)$-matrix $M_{\mathfrak{u}}$ whose entries are linear combinations of the components of an $m$-tuple of real parameters, where $m$ is the dimension of $\mathfrak{u}$ as a vector space, and where each realization of $M_{\mathfrak{u}}$ as a real matrix is understood to result from replacing its $m$ parameters with the members of any set of $m$ members of $\mathbb{R}$.
@@
For an $n$-dimensional algebra with elements $s=(x_{1},x_{2},\dots,x_{n})$ with respect to the standard basis, one initially takes $M_{\mathfrak{u}}$ to be a symmetric $(n\times n)$-matrix $M$ whose upper triangular entries are each a different member of a set of $\frac{n(n+1)}{2}$ real parameters. Computation of $s^{-1}$ produces a vector whose components are each a rational function of the $x_{1},x_{2},\dots,x_{n}$. So far, $Ms^{-1}$ can be understood as a parametrized $n$-dimensional vector field whose components (which can easily be computed) are dependent on the initial set of parameters in $M$ and the $x_{1},x_{2},\dots,x_{n}$. There are then $\frac{n(n-1)}{2}$ constraint equations that must be satisfied in order that (3.1) hold (for $L=M$). Each of these equations has the form,
@@
where $Ms^{-1}]_{k}$ indicates the $k$-th component of $Ms^{-1}$. As noted, we can write $s^{-1}=(\hat{x}_{1},\dots,\hat{x}_{n})$, where each $\hat{x}_{k}$ is a rational function of the components of $s$. Consequently, $Ms^{-1}]_{k}$ is a linear combination of rational functions of the components $\hat{x}_{1},\dots,\hat{x}_{n}$, where the coefficients defining this linear combination are linear combinations of various of the above parameters. It follows that $\frac{\partial}{\partial x_{p}}([Ms^{-1}]_{q})$ is again a sum rational functions of the components $x_{1},\dots,x_{n}$, where this expression is again linear in the parameters. Thus, multiplying both sides of (3.6) by a polynomial determined by the products of the polynomial denominators of the latter rational functions, the resulting left-hand-side of (3.6) becomes a polynomial in the components of $s$, while remaining an expression that is linear in the parameters - and the resulting right-hand-side of (3.6) remains zero. Hence, satisfaction of each of the $\frac{n(n-1)}{2}$ constraint equations (3.6) ultimately reduces to the requirement that a particular multivariate polynomial be identically zero. After simplification of the polynomial to be a linear combination of independent monomials, each monomial term is then comprised of a coefficient consisting of a linear combination of some of the above $\frac{n(n+1)}{2}$ parameters, multiplying an expression consisting of a product of some of the components of $s$, each raised to some power. Therefore, each of the above coefficients of these (independent) monomials are required to be zero. Accordingly, the number of surviving independent parameters (i.e., parameters not required to be equated to some linear combination of other parameters, or set to 0) is in general reduced as the $\frac{n(n-1)}{2}$ constraint equations are successively treated. $M$ is then modified as necessary, as each successive constraint equation is treated. After treatment of all the constraint equations via the above steps defined by treatment of the constraint equations in some sequence, the final number $m$ of surviving independent parameters is the dimension of $\mathfrak{u}$, and the final entries of $M$ evidently yield the requisite $M_{\mathfrak{u}}$. ∎
@@ 4.1. The transform of an uncurling metric
Let $L$ be an uncurling metric for some algebra $A$. For $s$ in an open ball centered at the multiplicative identity $\mathbf{1}$ and devoid of non-units, with $\|{\bf 1}\|^{2}\equiv\mathbf{1}\cdot\mathbf{1}$, the associated unital norm $\ell(s)$ is the function
@@
The set of unital norms associated with an algebra, which is an abelian group $\mathfrak{u}^{d}_{A}$ with respect to the usual point-wise operation of multiplication of functions, is called the unital norm group, also known as the dual of the anti-rotor. $\mathfrak{u}^{d}_{A}$ may also be denoted as $\mathfrak{u}^{d}$ if no confusion arises.
@@
The duality referred to in Definition 4.1 is easily appreciated in its relation to integral transforms. Equation (4.2) can be rewritten as $\log\ell(s)=\frac{1}{\|\mathbf{1}\|^{2}}\int_{{\bf 1}}^{s}\left[Lt^{-1}\right]\cdot\mathbf{d}t$. We can view this as a (forward) integral transform, mapping $Ls^{-1}$ to $\log\ell(s)$. But, based on the assumed standard basis, an expression for $s^{-1}$ is known for any given algebra (i.e., matrix $L$ can be extracted from $Ls^{-1}$, and $Ls^{-1}$ can be calculated from $L$), so this effectively results in a “forward transform” of $L$ to $\ell(s)$. Equation (4.2) also implies (4.1), the latter being an (inverse) integral transform from $\log\ell(s)$ to $Ls^{-1}$ in the sense of distributions. Thus we have an “inverse transform” of $\ell(s)$ to $L$. This duality is specifically defined with respect to the multiplicative inversion operation $s\mapsto s^{-1}$.

### v37

@@ 1. Introduction
As the means of understanding algebra structure, identification of the invariants of an algebra is a task of fundamental importance as well as an open-ended endeavor given that the algebra classification problem is apparently intractable [1, 2]. While classical algebra isomorphism invariants include things like an algebra’s dimension, types of ideals, and invariants stemming from the Jordan-Hölder Theorem, more recently they have derived from module categories, e.g., involving homological formalisms such as found in [3, 4, 5]. The invariants we present here belong to a different class.
@@ 2. The Skewer and the anti-rotor
In (2.1), $\text{Sym}_{n}$ is the linear space of all possible metric tensors relevant to $\mathbb{R}^{n}$ (inclusive of metrics that are not necessarily positive-definite and not necessarily nondegenerate), and $\delta\text{Rot}_{n}$ is the linear space of all possible infinitesimal rotation tensors relevant to $\mathbb{R}^{n}$. Equation (2.2) is interpreted as follows. An algebra element is understood to be a column vector $s=(x_{1},\dots,x_{n})^{T}$, we define $\mathbf{d}s\equiv(dx_{1},\dots,dx_{n})^{T}$, and $\mu\in\text{Sym}_{n}$ is thereby a real symmetric $(n\times n)$-matrix. On the right-hand-side of (2.2), $d$ represents the exterior derivative operation, thus acting on the 1-form $(f(s))^{T}\mu\,\mathbf{d}s=\big(\mu f(s)\big)\cdot\mathbf{d}s$. Consequently, the right-hand-side of (2.2) yields the infinitesimal rotation tensor associated with the vector field $f(s)$ with respect to the metric $\mu$ - or, equivalently, the infinitesimal rotation tensor of vector field $\mu f(s)$ with respect to the Euclidean metric.
@@
$\not{\text{S}}_{f}$ is a linear transformation mapping symmetric tensors to skew-symmetric tensors, so we refer to $\not{\text{S}}_{f}$ as a “Skewer” (the slash through “symmetric” S represents a skewer, as with a shish-kabob). Since we assume the standard basis, $\text{Sym}_{n}$ can be represented by a parametrized symmetric $(n\times n)$-matrix $M$ with $\frac{n(n+1)}{2}$ independent parameters, i.e., it is given as a symmetric matrix with a different real parameter in each of its upper triangular entries. Every possible metric tensor is given by some realization of the set of parameters (a realization being the result of any particular choice of $\frac{n(n+1)}{2}$ real values for the members of the set of parameters). Consequently, $M$ is mapped by a Skewer to a parametrized skew-symmetric matrix representing the resulting infinitesimal rotation tensors. Each entry not on the main diagonal of the latter matrix is a rational function of the $n$ components of the elements of algebra $A$, and the above $\frac{n(n+1)}{2}$ parameters appear as coefficients of constituent rational function terms which sum to the entry.
While the image of a Skewer is thereby complicated, the kernel of a Skewer is much less so for one and possibly two reasons. First, the components of the elements of $A$ do not appear. Second, there is a lot of computational evidence that kernel of the Skewer is independent of the choice of $f$ over the class of functions that are dependent on the algebra’s product operation. Note that for $f$ not dependent on the algebra product (i.e., of the form $f(s)=\alpha s^{0}+\beta s^{1}$, $\alpha,\beta\in\mathbb{R}$), the kernel of the Skewer is simply $\text{Sym}_{n}$ itself.
@@
We will henceforth adopt “anti-rotor” as a more descriptive term for the kernel of the Skewer, because each of the latter’s members annihilates the curl of the relevant vector fields. For definiteness, in the first seven sections of this paper we will work with $f(s)=s^{-1}$, since this choice is associated with a productive “dual” of the anti-rotor. The issue of the anti-rotor’s independence from the algebra-dependent vector fields is considered in Section 8.
@@ 3. The anti-rotor and resulting invariants
where we have used Definition 3.1, the commutativity of a differential with a linear transformation, and $(Ks_{1})^{-1}=K(s_{1}^{-1})$ (since $K$ is an isomorphism). It follows from (3.3) that every uncurling metric of algebra $A_{2}$ is associated with an uncurling metric of $A_{1}$, i.e., from the last equality in (3.3) we have $K^{T}L_{2}K\in\mathfrak{u}_{A_{1}}$. The implied map from $\mathfrak{u}_{A_{2}}$ to $\mathfrak{u}_{A_{1}}$ is obviously injective since $K^{T}L_{a}K=K^{T}L_{b}K$ implies $L_{a}=L_{b}$ due to the invertibility of $K$. The isomorphism hypothesis allows us to apply an analogous argument employing $K^{-1}$, indicating that for $L_{1}\in\mathfrak{u}_{A_{1}}$ we have the associated $(K^{-1})^{T}L_{1}K^{-1}\in\mathfrak{u}_{A_{2}}$, implying an injective map from $\mathfrak{u}_{A_{1}}$ to $\mathfrak{u}_{A_{2}}$. The theorem then follows from the bijection implied by the Cantor-Schoeder-Bernstein Theorem. ∎
@@
For invertible elements $s=(s_{ij})\in M_{n}(\mathbb{R})$, with det(s)(s) being the determinant of $s$, we have the well-known,
@@
where Adj(s)(s) is the adjugate matrix, and the left-hand-side is considered to be a matrix according to $\nabla\text{det}(s)]\equiv\left(\frac{\partial\,\text{det}(s)}{\partial s_{ij}}\right)$. The first of these equalities is a special case of Jacobi’s formula, and the second follows from the classical matrix inverse expression resulting from Laplace expansion. But equivalently, (3.4) can be understood as equating three vectors, where the matrix format of each bracketed expression is identified with the vector given by the ordered $n^{2}$-tuple obtained by sequentially appending its $n$ columns, and the transpose operation is now understood as a particular involutive $(n^{2}\times n^{2})$-matrix, which we will denote as $\mathcal{T}$. In this vector format, after dividing both sides by $\text{det}(s)$, (3.4) becomes
@@
Now suppose the left regular representation of $\bar{A}$ is not self-adjoint as a subalgebra of $M_{n}(\mathbb{R})$. Then there exists an epimorphism $\Omega:\bar{A}\rightarrow\underline{A}$ such that $\underline{A}$ is a simple algebra, and therefore $\underline{A}$ has an uncurling metric $\underline{L}$. A simple modification of the first three sentences of the proof of Theorem 3.2 (where we exchange isomorphism $K$ for epimorphism $\Omega$), indicates that $\Omega^{T}\underline{L}\Omega$ satisfies (3.1) and is thus an uncurling metric for $\bar{A}$ - along with all its scalar multiples. That is, with respect to the elements $\underline{s}\in\underline{A}$ and the elements $\bar{s}\in\bar{A}$, we consider the linear function $\underline{s}=\Omega\bar{s}$. Then, $0=d\left((\underline{s}^{-1})^{T}\underline{L}\mathbf{d}\underline{s}\right)=d\left(((\Omega\bar{s})^{-1})^{T}\underline{L}\mathbf{d}(\Omega\bar{s})\right)=d\left((\bar{s}^{-1})^{T}\big[\Omega^{T}\underline{L}\Omega\big]\mathbf{d}\bar{s}\right)$, where we have used Definition 3.1, the commutativity of a differential with a linear transformation, and $(\Omega\bar{s})^{-1}=\Omega(\bar{s}^{-1})$ (since $\Omega$ is an epimorphism of rings). ∎
@@
The members of $\mathfrak{u}$ are the realizations of a symmetric $(n\times n)$-matrix $M_{\mathfrak{u}}$ whose entries are linear combinations of the components of an $m$-tuple of real parameters, where $m$ is the dimension of $\mathfrak{u}$ as a vector space, and where a realization of $M_{\mathfrak{u}}$ as a real matrix is understood to result from replacing its $m$ parameters with the members of any set of $m$ members of $\mathbb{R}$.
@@
For an $n$-dimensional algebra with elements $s=(x_{1},x_{2},\dots,x_{n})$ with respect to the standard basis, one takes an initial candidate for $M_{\mathfrak{u}}$ to be the symmetric $(n\times n)$-matrix $M$ whose upper triangular entries are each a different member of a set of $\frac{n(n+1)}{2}$ real parameters (in effect, the initial candidate is taken to be all of $\text{Sym}_{n}$). Computation of $s^{-1}$ produces a vector whose components are each a rational function of the $x_{1},x_{2},\dots,x_{n}$. So far, $Ms^{-1}$ can be understood as a parametrized $n$-dimensional vector field whose components (which can easily be computed) are dependent on the initial set of parameters in $M$ and the $x_{1},x_{2},\dots,x_{n}$. Now, there are $\frac{n(n-1)}{2}$ constraint equations that must be satisfied in order that (3.1) hold for a candidate $L=M$. Each of these equations has the form,
@@
where $Ms^{-1}]_{k}$ indicates the $k$-th component of $Ms^{-1}$. It is easily recognized that in general $M$ will need to be modified in order to satisfy these constraint equations, and this is accomplished as follows.
As noted, we can write $s^{-1}=(\hat{x}_{1},\dots,\hat{x}_{n})$, where each $\hat{x}_{k}$ is a rational function of the components of $s$. Consequently, $Ms^{-1}]_{k}$ is a linear combination of rational functions of the components $\hat{x}_{1},\dots,\hat{x}_{n}$, where the coefficients defining this linear combination are linear combinations of various of the above parameters. It follows that $\frac{\partial}{\partial x_{p}}([Ms^{-1}]_{q})$ is again a sum rational functions of the components $x_{1},\dots,x_{n}$, where this expression is again linear in the parameters. Thus, multiplying both sides of (3.6) by a polynomial determined by the polynomial denominators of the constituent rational functions, the left-hand-side of (3.6) can be converted into a polynomial in the components of $s$, while remaining an expression that is linear in the parameters - and the right-hand-side of (3.6) remains zero under this procedure. In this way, satisfaction of each of the $\frac{n(n-1)}{2}$ constraint equations (3.6) ultimately reduces to the requirement that a particular multivariate polynomial be identically zero. After simplification of the polynomial to be a linear combination of independent monomials, each monomial term is then comprised of a coefficient consisting of a linear combination of some of the above $\frac{n(n+1)}{2}$ parameters, multiplying an expression consisting of a product of some of the components of $s$, each raised to some power. Therefore, each of the above coefficients of these (independent) monomials are required to be zero. Accordingly, one treats each of the $\frac{n(n-1)}{2}$ constraint equations in the above fashion one after the other in some sequence of steps. After each individual step, the number of surviving independent parameters (i.e., parameters not required to be equated to some linear combination of other parameters, or set to 0) is in general reduced as the $\frac{n(n-1)}{2}$ constraint equations are successively treated. $M$ is thereby sequentially modified as necessary as each successive constraint equation is treated. After treatment of all the constraint equations by executing the above steps in some sequence, the final number $m$ of surviving independent parameters is the dimension of $\mathfrak{u}$, and the final entries of the sequentially modified $M$ evidently yield the requisite $M_{\mathfrak{u}}$. ∎
@@ 4.1. The transform of an uncurling metric
Let $L$ be an uncurling metric for some algebra $A$. For $s$ in an open ball $D_{A}$ centered at the multiplicative identity $\mathbf{1}$ and devoid of non-units, with $\|{\bf 1}\|^{2}\equiv\mathbf{1}\cdot\mathbf{1}$, the associated unital norm $\ell:D_{A}\rightarrow\mathbb{R}$ is the function determined by,
@@
The set of unital norms associated with an algebra, which is an abelian group $\mathfrak{u}^{d}_{A}$ with respect to the usual point-wise operation of multiplication of real functions, is called the unital norm group, also known as the dual of the anti-rotor. $\mathfrak{u}^{d}_{A}$ may also be denoted as $\mathfrak{u}^{d}$ if no confusion arises.
@@
The duality referred to in Definition 4.1 is easily appreciated in its relation to integral transforms. Equation (4.2) can be rewritten as $\log\ell(s)=\frac{1}{\|\mathbf{1}\|^{2}}\int_{{\bf 1}}^{s}\left[Lt^{-1}\right]\cdot\mathbf{d}t$. We can view this as a (forward) integral transform, mapping $Ls^{-1}$ to $\log\ell(s)$. But, based on the assumed standard basis, an expression for $s^{-1}$ is known for any given algebra (i.e., matrix $L$ can be extracted from $Ls^{-1}$, and $Ls^{-1}$ can be calculated from $L$), so this effectively results in a “forward transform” of $L$ to $\ell(s)$ (of course, this transform is restricted to the space of uncurling metrics). Equation (4.2) also implies (4.1), the latter being an (inverse) integral transform from $\log\ell(s)$ to $Ls^{-1}$ in the sense of distributions. Thus we have an “inverse transform” of $\ell(s)$ to $L$ (restricted to the space of unital norms). This duality is specifically defined with respect to the multiplicative inversion operation $s\mapsto s^{-1}$.